OpenStudy (whalenjacob):

What is the solution to the following equation? 5(2x − 6) + 20 = 10 9 5 3 2

OpenStudy (whalenjacob):

@mathmale

OpenStudy (kittiwitti1):

do you know what the distributive property is

OpenStudy (whalenjacob):

sorta

OpenStudy (whalenjacob):

im going to try and solve it and you tell me if i did it right

OpenStudy (kittiwitti1):

okay

OpenStudy (whalenjacob):

i got 5?

OpenStudy (whalenjacob):

idk

OpenStudy (kittiwitti1):

sorry trying to multitask 2 questions and my own homework

OpenStudy (whalenjacob):

sorry i got 2

OpenStudy (whalenjacob):

is that right?

OpenStudy (raisengupta):

2 is correct

OpenStudy (whalenjacob):

ok thankyou so much

OpenStudy (whalenjacob):

i have another

OpenStudy (whalenjacob):

What is the value of z in the equation 4(2z + 3) = 12?

OpenStudy (whalenjacob):

−13 12 0 12

OpenStudy (kittiwitti1):

\[5(2x-6)+20=10\]\[(5\times2x)-(5\times6)=-10\]\[10x-30=-10\]\[10x=30-10\]\[10x=20\]yeah you got it right sorry for taking so long to type

OpenStudy (whalenjacob):

hang on im almost done

OpenStudy (whalenjacob):

i got 0? ithink that is right

OpenStudy (kittiwitti1):

(oops forgot one)\[x=\frac{20}{10}=2\] okay next one:\[4(2z+3)=12\]\[(4\times2x)+(4\times3)=12\]\[8x+12=12\]\[8x=0\]\[\frac{0}{8}=x=0\]yep

OpenStudy (whalenjacob):

ok

OpenStudy (whalenjacob):

How many solutions does the equation y + 12 = y + 10 + 2 have? Zero One Two Infinitely many

OpenStudy (kittiwitti1):

10+2 = 12 both sides are equal. what do you think?

OpenStudy (whalenjacob):

i think its infinite many solutions

OpenStudy (kittiwitti1):

\(\text{Correct}\) ☺

OpenStudy (whalenjacob):

ok

OpenStudy (whalenjacob):

What is the value of y in the equation 4(5y − 8 − 2) = 185 − 15? 8.5 9.7 10.5 42.5

OpenStudy (whalenjacob):

10.5?

OpenStudy (kittiwitti1):

\[4(5y-8-2)=185-15\]\[4(5y-10)=170\]\[(4\times5y)-(4\times10)=170\]\[20y-40=170\]\[20y=130\]\(y=\frac{130}{20}=\color{red}{6.5}\)

OpenStudy (kittiwitti1):

wait no I messed up lol

OpenStudy (kittiwitti1):

okay starting from \(20y-40=170\) again...\[20y=210\]\[y=\frac{200}{20}=10.5\]yeah you got it right

OpenStudy (kittiwitti1):

210* not 200 sorry

OpenStudy (whalenjacob):

ok

OpenStudy (whalenjacob):

this is the last one i dont get it

OpenStudy (whalenjacob):

The steps below show the incomplete solution to find the value of x for the equation 6x + 10 − 4x = −4 − 11: Step 1: 6x + 10 − 4x = −4 − 11 Step 2: 6x + 10 − 4x = −15 Step 3: 2x + 10 = −15 Which of these is most likely the next step? 2x = 5 2x = −25 2x = 25 2x = 150

OpenStudy (kittiwitti1):

they gave you the first 3 steps to solving the equation \(6x+10-4x=-4-11\) they want you to continue

OpenStudy (whalenjacob):

ok

OpenStudy (kittiwitti1):

so now you have \(2x+10=-15\) how will you find x from this?

OpenStudy (whalenjacob):

is it 2x=5?

OpenStudy (kittiwitti1):

no, close when you move the 10 over to the right side it becomes -10\[2x\cancel{+10-10}=-15-10\]\[2x=-25\]

OpenStudy (whalenjacob):

2x = −25

OpenStudy (whalenjacob):

thank you much

OpenStudy (kittiwitti1):

no problem. happy studying ☺

OpenStudy (whalenjacob):

i gave you a medal

OpenStudy (kittiwitti1):

I saw xD

OpenStudy (kittiwitti1):

don't forget to close this problem P: